Two cars leave Charlotte at the same time and travel in opposite directions along I- 77. One car travels 14 miles per hour faster than the other car. After four hours they are 480 miles apart. What is the speed of the slower car?
Answer by checkley75(3666)

4x+4(x+14)=480
8x+56=480
8x=480-56
8x=424
x=424/8
x=53 mph. for the slower car.
Proof:
4*53+4(53+14)=480
212+4*67=480
212+268=480
480=480